# How to Train Your LLM Web Agent: A Statistical Diagnosis

### v1

@@ 6 Ablation and Sensitivity Analysis
Figure 4: Bootstrap analysis ($n=1000$ samples) of hyperparameter optimization across different SFT compute budgets on training held out tasks. Each subplot examines a different hyperparameter, including increasing SFT compute: the base instruct model (left), +2.5×$\times 10^{18}$ SFT FLOPs (middle), and +7.6e×$\times 10^{18}$ SFT FLOPs (right). For each hyperparameter-compute combination, the top panel shows relative reward performance with error bars indicating 95% confidence intervals, while the bottom panel displays win rates representing the percentage of bootstrap iterations where each parameter value achieved maximum performance. Results demonstrate that optimal hyperparameter values shift as model pre-training compute increases, suggesting that hyperparameter selection should be adapted to the computational budget allocated to SFT.
@@
We simulate re-running hyperparameter configurations and selecting the best-performing ones using the method described in Section 3.1. This is done across three checkpoints: the base LLaMA 3.1 8B Instruct model and two warm-started variants with an additional $2.5\times 10^{18}$ and $7.6\times 10^{18}$ FLOPs of supervised fine-tuning, respectively, to assess variations across compute budgets. We evaluate the held-out goals and held-out tasks performance of 10 hyper-parameters across 1,370 runs. We report final held-out task performance to verify generalization in Appendix D finding no significant deviations between held-out tasks and held-out goals parameters.
@@
Figure 4 displays our findings, which we summarize as follows. Curriculum learning is beneficial when starting RL from scratch but becomes detrimental after warm-starting, likely because warm-started models already perform well on easy tasks, and curriculum forces them to overfocus on harder ones. Error log feedback helps when there’s no SFT but otherwise does not, likely because SFT warmup removes many common errors made by weaker models. A decoding temperature of 0.25 consistently yields the best results, striking a balance between exploration and exploitation; lower values led to under-exploration and were discarded. Grouped-relative advantage helps only after SFT, while using raw rewards works better when starting directly from the Instruct model, possibly due to how advantage scaling interacts with the initial weights. Zero-advantage filtering improves training across most settings by ensuring batches focus on informative updates. Standard-deviation normalized advantage, as noted by [14], seems to aid performance under less finetuned models and decrease in value the more finetuning is done. Importance ratio correction and trust region, though standard, also hurt models with little or no SFT, likely because conservative updates slow down learning. In contrast, for models that start from a stronger SFT checkpoint, these mechanisms can help stabilize training and avoid catastrophic updates. For the learning rate, the larger value (1e-16) generally worked better. Effective batch size of 512 appears to be generally be a safe and robust choice in all our experiments. Finally, regarding the discount rate, values above 0.9 work well for models with little or no SFT warmup, while heavily warm-started models benefit from a lower rate around 0.5—likely because it encourages the agent to optimize more aggressively on tasks it already handles well.
@@ 7 Related Work
Recent advancements have sought to bridge the gap in training LLM agents for multi-step environments, with approaches like WebRL [16] and SWEET-RL [26] demonstrating significant progress. WebRL employs a self-evolving curriculum to address the challenges of sparse feedback and task scarcity, successfully enhancing the performance of open LLMs in web-based tasks [16]. Similarly, SWEET-RL introduces a hierarchical structure that enables effective credit assignment over multiple turns, improving policy learning and generalization in collaborative reasoning tasks [26]. These studies collectively illustrate the necessity of adapting RL techniques to accommodate the complexities of multi-step interactions, paving the way for more capable and versatile LLM agents. Similar in spirit to our work, [4] propose an empiral study on the inference cost of trained LLM web agents. An extended version of the related work is provided in Appendix C.
@@
Recent work has begun closing the gap for LLM agents that must reason over multiple steps. WebRL [16] introduces a self-evolving curriculum that tackles sparse feedback and task scarcity, substantially boosting open-source agents on web tasks. SWEET-RL [26] adds a hierarchical credit-assignment scheme spanning many turns, improving both learning stability and generalization in collaborative reasoning. Together, these studies underscore the need to adapt RL techniques to the intricacies of long-horizon interaction, paving the way for more capable, versatile agents. Similar in spirit, [4] provide an empirical analysis of inference costs for trained LLM web agents. An extended discussion of related work appears in Appendix C.
@@ Conclusion.
We present a statistically grounded study on training LLM web agents, analyzing the trade-off between SFT and RL. We perform a random sweep across 1,370 configurations to identify optimal hyperparameter choices which—importantly— we find can vary across compute budgets. Using the bootstrap prescribed hyper-parameters we show that branching into RL early—but not immediately—after SFT achieves better performance–compute trade-offs, matching peak SFT with 45% less compute and closing the gap with closed-source agents. Our findings offer a reproducible, budget-aware blueprint for advancing open-source LLM web agents in complex multi-step environments.
Appendix A Extended Learning and Saturation Analysis
@@
Appendix B Deriving Compute Cost
@@
$\text{FLOPs}_{\text{attn}}=12\times\text{(number of layers)}\times\text{(hidden size)}^{2}\\ \times\left(1+\frac{\text{number of query groups}}{\text{number of attention heads}}+\frac{\text{sequence length}}{\text{hidden size}}\right)$ (6)
@@
[…]
@@
Appendix C Extended Related Work
@@
Best prectices in deep RL.
@@
Appendix D Test Set Hyper-Parameter Bootstrap Analysis
@@
Appendix E Random Search Space
@@
Appendix F Compute Allocation Hyperparameter Selection
For consistency, we select a single set of hyperparameters for all runs in Figure 1. To do so, we use the bootstrap algorithm described in Algorithm 1 where we use all the runs over all SFT checkpoints as input into Results. We then analyze the results of this bootstrap and use them to obtain the reported results.

### v3

@@ 6 Ablation and Sensitivity Analysis
[…]
@@
We simulate re-running hyperparameter configurations and selecting the best-performing ones using the method described in Section˜3.1. This is done across three checkpoints: the base LLaMA 3.1 8B Instruct model and two warm-started variants with an additional $2.5\times 10^{18}$ and $7.6\times 10^{18}$ FLOPs of supervised fine-tuning, respectively, to assess variations across compute budgets. We evaluate the held-out goals and held-out tasks performance of 10 hyper-parameters across 1,370 runs. We report final held-out task performance to verify generalization in Appendix˜G finding no significant deviations between held-out tasks and held-out goals parameters.
@@
Figure˜3 displays our findings, which we summarize as follows. Curriculum learning is beneficial when starting RL from scratch but becomes detrimental after warm-starting, likely because warm-started models already perform well on easy tasks, and curriculum forces them to overfocus on harder ones. Error log feedback helps when there’s no SFT but otherwise does not, likely because SFT warmup removes many common errors made by weaker models. A decoding temperature of 0.25 consistently yields the best results, striking a balance between exploration and exploitation; lower values led to under-exploration and were discarded. Grouped-relative advantage helps only after SFT, while using raw rewards works better when starting directly from the Instruct model, possibly due to how advantage scaling interacts with the initial weights. Zero-advantage filtering improves training across most settings by ensuring batches focus on informative updates. Standard-deviation normalized advantage, as noted by [17], seems to aid performance under less finetuned models and decrease in value the more finetuning is done. Importance ratio correction and trust region, though standard, also hurt models with little or no SFT, likely because conservative updates slow down learning. In contrast, for models that start from a stronger SFT checkpoint, these mechanisms can help stabilize training and avoid catastrophic updates. For the learning rate, the larger value (1e-6) generally worked better. Effective batch size of 512 appears to be generally be a safe and robust choice in all our experiments. Finally, regarding the discount rate, values above 0.9 work well for models with little or no SFT warmup, while heavily warm-started models benefit from a lower rate around 0.5 likely because it encourages the agent to optimize more aggressively on tasks it already handles well. A further hypothesis is that the large deviation in optimal hyperparameters between the base and warm-start models could be attributed to the significant entropy reduction introduced by SFT (see Appendix˜H).
@@ 7 Related Work
Recent advancements have sought to bridge the gap in training LLM agents for multi-step environments, with approaches like WebRL [19] and SWEET-RL [30] demonstrating significant progress. WebRL employs a self-evolving curriculum to address the challenges of sparse feedback and task scarcity, successfully enhancing the performance of open LLMs in web-based tasks [19]. Similarly, SWEET-RL introduces a hierarchical structure that enables effective credit assignment over multiple turns, improving policy learning and generalization in collaborative reasoning tasks [30]. These studies collectively illustrate the necessity of adapting RL techniques to accommodate the complexities of multi-step interactions, paving the way for more capable and versatile LLM agents. Similar in spirit to our work, [4] propose an empiral study on the inference cost of trained LLM web agents. An extended version of the related work is provided in Appendix˜F.
@@
Recent work has begun closing the gap for LLM agents that must reason over multiple steps. WebRL [19] introduces a self-evolving curriculum that tackles sparse feedback and task scarcity, substantially boosting open-source agents on web tasks. SWEET-RL [30] adds a hierarchical credit-assignment scheme spanning many turns, improving both learning stability and generalization in collaborative reasoning. Together, these studies underscore the need to adapt RL techniques to the intricacies of long-horizon interaction, paving the way for more capable, versatile agents. Similar in spirit, [4] provide an empirical analysis of inference costs for trained LLM web agents. An extended discussion of related work appears in Appendix˜F.
@@ Conclusion.
We present a statistically grounded study on training LLM web agents, analyzing the trade-off between SFT and RL. We perform a random sweep across 1,370 configurations to identify optimal hyperparameter choices, finding interesting variation across compute budgets. Using the bootstrap prescribed hyper-parameters we show that branching into RL early, but not immediately, after SFT achieves better performance–compute trade-offs, matching peak SFT with 45% less compute and closing the gap with closed-source agents. Our findings offer a reproducible, budget-aware blueprint for advancing open-source LLM web agents in complex multi-step environments.
Appendix A Compute-peformance frontier for Qwen2.5 7B
Figure 4: Compute–performance frontier on MiniWoB++ (results averaged over two seeds) for Qwen2.5 7B. The blue curve shows pure SFT on teacher demonstrations. Warm-colored curves represent hybrid runs that branch off from SFT checkpoints and continue training with RL. Early transitions to RL push the Pareto frontier achieving higher success rates for the same compute and is the only approach able to achieve over $30\%$ improvement on both held-out goals (left) and held-out tasks (right).
Appendix B Compute Infrastructure
Appendix C Extended Learning and Saturation Analysis
@@
Appendix D Deriving Compute Cost
@@
$\text{FLOPs}_{\text{MLP}}$ $=18\times\text{(number of layers)}\times\text{(hidden size)}\times\text{FFN}$ (8)
$\text{FLOPs}_{\text{embed}}$ $=6\times\text{vocabulary size}\times\text{(hidden size)}$ (9)
@@
Appendix E Benchmark Descriptions
Appendix F Extended Related Work
@@
Best practices in deep RL.
@@
Appendix G Test Set Hyper-Parameter Bootstrap Analysis
@@
Appendix H Entropy
Figure˜9 reports log-entropy as a function of training compute (FLOPs) across the training run.
Figure 9: relationship between log-entropy and compute allocated to SFT warm-starting. We observe an initial sharp drop in log-entropy, followed by a period of stabilization and gradual decline. We speculate that this reduction in entropy, which makes rollouts more deterministic, may explain the difference in optimal hyperparameters between the base and warm-started checkpoints.
Appendix I Random Search Space
@@
Appendix J Compute Allocation Hyperparameter Selection
For consistency, we select a single set of hyperparameters for all runs in Figure˜1. To do so, we use the bootstrap algorithm described in Algorithm˜1 where we use all the runs over all SFT checkpoints as input into Results. We then analyze the results of this bootstrap and use them to obtain the reported results.
@@
Appendix K Correctness of the Bootsrap
One of the main motivations for using the boostrap algorithm for our hyper-parameter evaluation is the vast amount of literature showing its desrable properties for estimating the sampling distribution of data statistics.
The most important attribute for the robustness of the boostrap is its ability to exploit the Central Limit Theorem (CLT) to approximate the sampling distribution of a statistic. Using the CLT it can allow us to estimate its expectation and variability without parametric assumptions [12].
Specifically, in our use case, we want to evaluate, for each hyperparameter value $h$, how well it performs when paired with the best possible values of the other hyperparameters. The bootstrap algorithm yields an unbiased estimate for a target statistic given an initial $n$ samples (full runs in our case). Thus, we choose the statistic of interest to be the evaluation score $M$ conditioned on the hyper-parameter of interst $h$ assuming all other hyper-parameters $g$ are optimal:
$T_{h}(D)=\max\left\{M(h^{\prime},g):(h^{\prime},g)\in D,\ h^{\prime}=h\right\}$ (16)
Bootstrapping this yields an empirical distribution over $T_{h}$ values, providing an unbiased estimate and confidence intervals for the maximum score achievable for $h$ when other hyperparameters are optimal.
Additionally, there exists results showing the rate of convergence and consistency of the bootstrap algorithm which we directly exploit. It is a well known result that as the number of training runs $n\to\infty$, the bootstrap estimate $\hat{T}_{h}$ of the maximum achievable score for hyperparameter value $h$ becomes consistent. That is:
$\hat{T}_{h}\xrightarrow{p}T_{h}=\sup\left\{M(h,g):g\in\mathcal{G}\right\}$ (17)
This follows from the law of large numbers [24], assuming sufficient coverage of other hyperparameters $\mathcal{G}$. The rate of convergence depends on the tail behavior of the conditional distribution $M\mid H=h$. Since our evaluation metric is bounded in $[0,1]$, the sub-Gaussian assumption holds automatically by Hoeffding’s Lemma [14], ensuring concentration of the estimate. Under these conditions, the expected bias decays as $\mathcal{O}(1/n_{h})$, where $n_{h}$ is the number of runs with $H=h$.
These are well-known results, which is why the bootstrap is an extremely useful and simple algorithm for estimating any statistic of interest [11, 12].
Appendix L Train and Test Splits Used in the Experiments
We evaluate generalization by training only on the train split and reporting performance on held-out tasks from the test split. For the held-out goals metric, we instantiate goal variations using seed ranges $[0,2]$ [8] for training tasks (3 goals per task) and $[0,9]$ [8] for test tasks (10 goals per task). Below we list the exact task identifiers used in our experiments for both benchmarks. Task names are the registry keys from the respective environments.
WorkArena (Train: 24 tasks, Test: 7 tasks)
Train:
  • workarena.servicenow.all-menu
  • workarena.servicenow.create-hardware-asset
  • workarena.servicenow.create-incident
  • workarena.servicenow.filter-asset-list
  • workarena.servicenow.filter-change-request-list
  • workarena.servicenow.filter-hardware-list
  • workarena.servicenow.filter-service-catalog-item-list
  • workarena.servicenow.filter-user-list
  • workarena.servicenow.knowledge-base-search
  • workarena.servicenow.order-apple-mac-book-pro15
  • workarena.servicenow.order-development-laptop-p-c
  • workarena.servicenow.order-ipad-mini
  • workarena.servicenow.order-ipad-pro
  • workarena.servicenow.order-sales-laptop
  • workarena.servicenow.sort-change-request-list
  • workarena.servicenow.sort-hardware-list
  • workarena.servicenow.sort-incident-list
  • workarena.servicenow.sort-service-catalog-item-list
  • workarena.servicenow.sort-user-list
  • workarena.servicenow.single-chart-value-retrieval
  • workarena.servicenow.create-change-request
  • workarena.servicenow.order-loaner-laptop
  • workarena.servicenow.order-standard-laptop
  • workarena.servicenow.create-problem
Test:
  • workarena.servicenow.create-user
  • workarena.servicenow.filter-incident-list
  • workarena.servicenow.sort-asset-list
  • workarena.servicenow.impersonation
  • workarena.servicenow.order-apple-watch
  • workarena.servicenow.order-developer-laptop
  • workarena.servicenow.single-chart-min-max-retrieval
MiniWoB++ (Train: 99 tasks, Test: 23 tasks)
Train:
  • miniwob.ascending-numbers
  • miniwob.bisect-angle
  • miniwob.book-flight
  • miniwob.choose-date
  • miniwob.choose-date-easy
  • miniwob.choose-date-medium
  • miniwob.choose-date-nodelay
  • miniwob.choose-list
  • miniwob.circle-center
  • miniwob.click-button-sequence
  • miniwob.click-checkboxes-soft
  • miniwob.click-checkboxes-transfer
  • miniwob.click-collapsible-2
  • miniwob.click-collapsible-2-nodelay
  • miniwob.click-collapsible-nodelay
  • miniwob.click-color
  • miniwob.click-dialog
  • miniwob.click-dialog-2
  • miniwob.click-link
  • miniwob.click-menu
  • miniwob.click-menu-2
  • miniwob.click-scroll-list
  • miniwob.click-shape
  • miniwob.click-tab
  • miniwob.click-tab-2
  • miniwob.click-tab-2-hard
  • miniwob.click-tab-2-medium
  • miniwob.click-test
  • miniwob.click-test-2
  • miniwob.click-test-transfer
  • miniwob.click-widget
  • miniwob.copy-paste
  • miniwob.copy-paste-2
  • miniwob.count-shape
  • miniwob.count-sides
  • miniwob.daily-calendar
  • miniwob.drag-box
  • miniwob.drag-circle
  • miniwob.drag-cube
  • miniwob.drag-items
  • miniwob.drag-items-grid
  • miniwob.drag-shapes
  • miniwob.drag-shapes-2
  • miniwob.drag-sort-numbers
  • miniwob.draw-circle
  • miniwob.draw-line
  • miniwob.email-inbox
  • miniwob.email-inbox-delete
  • miniwob.email-inbox-forward
  • miniwob.email-inbox-forward-nl
  • miniwob.email-inbox-forward-nl-turk
  • miniwob.email-inbox-important
  • miniwob.email-inbox-noscroll
  • miniwob.email-inbox-reply
  • miniwob.email-inbox-star-reply
  • miniwob.enter-date
  • miniwob.enter-text
  • miniwob.enter-text-dynamic
  • miniwob.enter-time
  • miniwob.find-greatest
  • miniwob.find-word
  • miniwob.focus-text-2
  • miniwob.form-sequence
  • miniwob.form-sequence-2
  • miniwob.generate-number
  • miniwob.grid-coordinate
  • miniwob.guess-number
  • miniwob.highlight-text
  • miniwob.hot-cold
  • miniwob.identify-shape
  • miniwob.login-user
  • miniwob.login-user-popup
  • miniwob.multi-layouts
  • miniwob.multi-orderings
  • miniwob.navigate-tree
  • miniwob.odd-or-even
  • miniwob.order-food
  • miniwob.phone-book
  • miniwob.read-table
  • miniwob.read-table-2
  • miniwob.resize-textarea
  • miniwob.right-angle
  • miniwob.scroll-text
  • miniwob.scroll-text-2
  • miniwob.search-engine
  • miniwob.sign-agreement
  • miniwob.simple-algebra
  • miniwob.social-media
  • miniwob.social-media-all
  • miniwob.social-media-some
  • miniwob.text-editor
  • miniwob.text-transform
  • miniwob.tic-tac-toe
  • miniwob.use-autocomplete
  • miniwob.use-autocomplete-nodelay
  • miniwob.use-colorwheel
  • miniwob.use-colorwheel-2
  • miniwob.use-spinner
  • miniwob.visual-addition
Test:
  • miniwob.buy-ticket
  • miniwob.click-button
  • miniwob.click-option
  • miniwob.click-pie-nodelay
  • miniwob.drag-single-shape
  • miniwob.email-inbox-nl-turk
  • miniwob.enter-text-2
  • miniwob.find-midpoint
  • miniwob.focus-text
  • miniwob.simple-arithmetic
  • miniwob.stock-market
  • miniwob.use-slider-2
  • miniwob.click-checkboxes
  • miniwob.click-checkboxes-large
  • miniwob.click-collapsible
  • miniwob.click-pie
  • miniwob.click-shades
  • miniwob.click-tab-2-easy
  • miniwob.enter-password
  • miniwob.form-sequence-3
  • miniwob.highlight-text-2
  • miniwob.unicode-test
  • miniwob.use-slider